It is 2:07 AM and I am fried. Please help.
Solve by the addition method
5x -3y= -20
3x + 2y= 7

Answer by uma(370)

5x - 3y = - 20 ---------------(1)
3x + 2y = 7 -----------------(20
(1) x 2 gives 10x - 6y = - 40
(2) x 3 gives 9x + 6y = 21
Adding the above we get..
19x = - 19
==> 19x/19 = - 19/19
==> x = - 1
Plugging in x = -1 in (2)...
3(-1) + 2y = 7
==> -3 + 2y = 7
==> 2y = 7 + 3 [adding 3 to both the sides of the equation]
==> 2y = 10
==> 2y/2 = 10/2
==> y = 5
Thus x = -1, y = 5

Good Luck!!!